OpenStudy (anonymous):

the integral of 1/(1+4x^2) SHOW WORK PLEASE

OpenStudy (anonymous):

start let (1/2)tantheta

OpenStudy (anonymous):

x=(1/2)tantheta dx=?

OpenStudy (anonymous):

(1/2)sec^2 theta

OpenStudy (anonymous):

ok then substitute in your problem by x=(1/2)tantheta and dx=(1/2)sec^2theta dtheta